Please help my son with this word problem. I have tried and cannot. I would really appreciate your help. Thanks so much!
Tom is 4 years older than Jerry. Nine years ago, Tom was 5 times as old as Jerry was then. How old is each now?

Answer by checkley71(8403)

T=J+4 & T-9=5[J-9] OR (J+4)-9=5(J-9) OR J+4-9=5J-9 OR J+4=5J OR 4=4J OR J=1
THEN T=5*1 OR T=5 AGES 9 YEARS AGE FOR J & T.
THEY ARE NOW 10 & 14 (JERRY & TOM)